OpenStudy (anonymous):

Need help! Got through finding the first derivative, but not too sure about my answer. Here is the question: When f(x) = (x-2)e^x A) Find the intervals of increase of decrease B) Find the local max and mins C) Find the intervals of concavity and inflection points D) Use the information from parts A-C to sketch the graph.

OpenStudy (anonymous):

For the first derivative I got: f'(x) = e^x (x-1)

OpenStudy (danjs):

yes

OpenStudy (anonymous):

D

OpenStudy (anonymous):

So for the critical values, for when f'(x) = 0 I got x=1. But for the other critical value when f'(X)= undefined... I tried to put the first derivative in a fraction form but Im not sure Its right.

OpenStudy (anonymous):

When the first derivative in fraction form, I got: e^x - (x-2)* (e^x) all over (e^-x)^2

OpenStudy (anonymous):

so when f'(x) is undefined, does x= 0? Im not sure...

OpenStudy (danjs):

sorry my mouse died

OpenStudy (danjs):

yes x=1 has a horizontal tangent point, test either side of it to see what the function is doing

OpenStudy (anonymous):

f'(x)=0 gives x=1

OpenStudy (danjs):

is f '(x) positive(increasing) or negative(decreasing) function on etiher side of x=1, test any point

OpenStudy (danjs):

Just for practice, here is the actual graph of f, to compare your answers to..

OpenStudy (danjs):

since x=1 is the only critical point here, you have two intervals to test for inc/dec (-inf, 1) and (1, +inf)

OpenStudy (anonymous):

Really? what about when f'(x) = undefined, does that give us any critical values?

OpenStudy (danjs):

where does that happen?

OpenStudy (anonymous):

so you would have to do the first derivative in a fraction form and then whatever is at the bottom of that fraction, you set that to zero and solve...and whatever you get from that is another critical value. But i'm not even sure that I put the first derivative in its right fraction form

OpenStudy (danjs):

f ' (x) = (x-1)e^x what did you do next

OpenStudy (anonymous):

So I actually got what you got, but then I did it all over again trying to put it into fraction form and this is what I got: e^x - (x-2)* (e^x) all over (e^-x)^2

OpenStudy (danjs):

Oh, do you mean the original function f(x)

OpenStudy (anonymous):

Yea from the original function, I tried to find the first derivative in a fraction form..and well I got e^x - (x-2)* (e^x) all over (e^-x)^2

OpenStudy (danjs):

f(x) = (x-2)e^x = xe^x - 2e^x f(x) = xe^x - 2e^x

OpenStudy (danjs):

i dont understand what you did

OpenStudy (danjs):

using the product rule here , no quotient

OpenStudy (anonymous):

Okay, so I basically did: f(x)= (x-2)/ e^-x and then I did the quotient rule on this to find the first derivative in fraction form But I assumed to use the quotient rule because I took the e^x down

OpenStudy (anonymous):

If I took e^x down to become e^-x, wouldnt I be able to use the quotient rule? That seems logical to me

OpenStudy (danjs):

yes, that gets messy with ln functions

OpenStudy (danjs):

but it will come out to x=1 the same

OpenStudy (danjs):

when you set it to zero

OpenStudy (anonymous):

I didnt knowthat when (e^-x)^2 =0 it solves out to be x=1

OpenStudy (danjs):

e^(-2x) never is zero

OpenStudy (anonymous):

So we get out nothing when f'(x)= undefined? Either way I got inc from: (1,infinity) and decreasing from: (negative infinity, 1).

OpenStudy (danjs):

there is no undefined value for the derivative

OpenStudy (anonymous):

Okay good. And for local max, i got none. But for local min I got x= 1

OpenStudy (anonymous):

Then to find the intervals of concavity and inflection points I know I have to find the second derivative and I got f''(x) = e^x (x) which basically gives me x=0

OpenStudy (anonymous):

And if I am right with the critical value... do I do a number line that has 1 and 0 and test out points between them

OpenStudy (anonymous):



OpenStudy (danjs):

yes so far so good

OpenStudy (danjs):

so far you have from first derivative...